# You May Speak Freely: Improving the Fine-Grained Visual Recognition Capabilities of Multimodal Large Language Models with Answer Extraction

### v1

@@ 1 Introduction
Zero-shot visual classification has been revitalized by the introduction of Multimodal Large Language Models (MLLMs) [20, 9, 1, 21], where one can formulate classification datasets as Visual Question Answering (VQA) tasks with Multiple Choice Questions (MCQs) and a common prompt. However, fine-grained visual classification (FGVC) poses a barrier to traditional MCQ problems due to two issues: (1) a vastly higher number of choices (often hundreds to thousands) and (2) the niche concepts that each choice represents. Recent evaluations have found that state-of-the-art instruction-tuned MLLMs (e.g. LLaVA 1.5 [20, 21], InstructBLIP [9], GPT-4V [1]) suffer dramatic drops in accuracy when asked to name fine-grained categories. For instance, LLaVA-1.5 (13B) achieves near-perfect accuracy on coarse categories (like “bird” vs “not bird”), but only  1–2% accuracy on fine-grained species labels in iNaturalist [15].
Figure 2: Seemingly innocuous instruction changes greatly affect MLLM performance. When testing on the gull subset of CUB200[40], replacing one word can change the predictions of a model (left). When rewriting an instruction in a semantically equivalent way, performance varies substantially for Qwen-2.5VL (right). Each distribution comprises 15 semantically equivalent ways of writing the same instruction. These cases highlight the need to test LLMs over various equivalent inputs to ensure robustness.
Using LLMs to extract choices from pre-existing text appears to be a potential solution to this problem. Constraining LLM’s output format to ease extraction can simplify scoring but might hinder the model’s reasoning [41, 35, 43], whereas allowing free-form explanations improves reasoning at the cost of harder extraction [27, 35]. Many approaches prompt a strong LLM [8, 18] such as GPT4 [2] to perform second-stage answer extraction. However, the works that propose LLMs as choice extractor are either focused only on language-only tasks [24, 44] or do not extend beyond simple lettering or numbering schemes [23]. Furthermore, no significant treatment of retrieval nor the robustness of these results under input prompt variation has been considered.
@@ 2 Related Work
The earliest works first prompted an LLM directly for a response then did regex parsing to extract choices [42], but it was shown that direct open-ended querying often underperforms on fine-grained tasks [15]. While generating the full probability the text of each choice sequence is as an option, it remains difficult to implement in practice due to its computational cost [31]. The most dominant choice probability generation rank the log probabilities assigned to the option IDs (e.g. ”A/B/C/D”) from the first token prediction of the model. This approach is widely adopted in many different benchmarks and LLM evaluation studies [14, 45, 34, 19, 33].
Parallel to this, constrained decoding techniques seek to guarantee consistent outputs by guiding the generation process during sampling by constricting the output vocabulary [5, 22]. There are many open-source tools that implement this approach [10, 13, 11, 46] and while these methods ensure model responses lie within the choice set, recent evaluations have shown that they decrease performance [41, 35]. Our work sits in opposition to these recent results, where we posit that constrained decoding is advantageous in the fine-grained setting where previous approaches like first token prediction fail.
@@
Recent works have shown a mismatch difference between text-based outputs and token-probability metrics of LLMs on language-only tasks, revealing that a model’s first-token likelihood often fails to correspond with its generated answer [24, 44, 43]. Evaluating a model by reading its full text answer instead of only the first-token or forced letter has been shown to improve robustness, reinforcing the value of answer extraction methods [43]. Many works use large proprietary models like GPT-4 [2] to extract answers from free-form text [23, 8, 18].
Some recent methods have addressed the problem of answer extraction directly. xFinder [47] propose a lightweight model fine-tuned specifically to map an LLM’s free-form answer to the correct choice, reporting that their method outperforms a GPT-4 based extractor struggled in terms of extraction accuracy. Similarly, SLOT [41] proposes a model-agnostic, lightweight “Structured LLM Output Transformer” that is fine-tuned to post-process any LLM’s free-form text into JSON that exactly matches an arbitrary schema, eliminating the need to modify or constrain the generating model itself. xVerify [7] is an efficient answer verifier for reasoning model evaluations that extracts final answers from complex, multi-step LLM outputs and performs robust equivalence checking—trained on the VAR dataset of multi-model, multi-round–annotated question–answer pairs and achieving over 95% accuracy and F1, demonstrating strong generalizability. Contrary to these approaches the answer extraction method proposed in this work requires no additional training. We argue that answer extraction is not a critical bottleneck when using constrained decoding for fine-grained visual classification and show that LLMs perform decent answer extraction out-of-the-box on popular benchmarks.
@@ 3.1 Simulating User Variation
Where the words ”species” and ”bird” are updated to suite the benchmark dataset. Given [templatei], we use o3-high [29] to generate 14 additional prompt variations:
@@
Table 1: Number of forward passes needed for different methods in the second-stage answer extraction of nlg2choice. ”Full Prob” refers to the probability of a full class name $p(\texttt{[cname]})$ under the LLM, where [cname] is a common name of a class. “Yes/No” refers to the common $|\mathcal{X}|\cdot|\mathcal{Y}|$ method of calculating probability over all the classes $y_{i}\in\mathcal{Y}$. “Truncated Prob” refers to the process outlined in Sec. $3.$ The percentages in the “Yes/No” and ”Truncated Prob” columns refer to the throughput increase over caculating the full sequence probabilities ”Full Prob.” Every setting for truncation is faster than “Yes/No” except for one benchmark - NABirds [38]. The fastest method for each benchmark is bolded.
@@ 4 Experiments
We conduct experiments over a benchmark of 7 FGVC datasets: CUB [40] (200 classes), Flowers 102 [28] (102 classes), Stanford Cars [17] (196 classes), FGVC Aircrafts [25] (100 classes) and Food101 [6] (101 classes), NABirds [38] (555 classes), and iNaturalist-Birds [39] (1486 classes). We choose these benchmarks because they are well-known for fine-grained visual classification and are traditionally difficult for zero-shot visual models to solve. iNaturalist-Birds is the validation split of the iNat2021 dataset where the the ”class” field of each row is equal to ”Aves,” resulting in a test size of $14960$. For each dataset we use the full test set and the training split remains unused.
Evaluation Metrics   For the classification setting we use accuracy averaged over 15 semantically equivalent prompts, whereas for the retrieval setting we use Mean Average Precision (mAP). Choice lists are the names of classes themselves rather than letters or numbers. As compared to previous works [36], we perform the full-way class prediction rather than subsetting the choices to a 4-way prediction. However, we provide an experiment evaluating this setting in Appendix B.
@@
Implementation Details   Constrained decoding is implemented using the Outlines [^2] https://github.com/dottxt-ai/outlines library [46] and classes for each setting are detailed in Appendix C. We start by generating 15 prompts described in Sec. 3.1, then replacing the key features like domain (”birds,” ”flowers,” etc.) and class types (”species,” ”variant,” etc.) for each dataset. This allows us to compare directly between questions for a given model or dataset, and we also test the performance of appending class lists and common CoT methods. For our experiments, we test only open-source models. Specifically, we test the Qwen-2.5VL [4], Llama-3.2 [12], and InternVL3 [49] architectures. For hyperparameters, we use the recommended generation parameters as defined by the HuggingFace card and take max new tokens to be 512. Experiments were ran in a heterogeneous environment of single-GPU nodes. All nodes required a GPU with at least 23GB VRAM, 16 CPU cores, and 64GB RAM.
@@ 5.1 Zero-Shot Classification Performance
In Tab. 2 we report the performance of nlg2choice across various architectures and datasets. All architectures enjoyed a bump in performance across individual datasets and on average (+17.46, +8.49, +6.87 across Qwen-2.5VL, Llama-3.2V, and Intern3), but model-specific differences can be clearly seen. Likewise, we find that removing explicit answer formatting from the model (changing from “nlg2choice” and “nlg2choiceopen”). We find the Aves subset of iNaturalist [39] to be the hardest among all the benchmarks, resulting in the smallest increase in performance across all models.
For model-specific differences, we find that Qwen-2.5VL performs the best on average using nlg2choice, having an average accuracy of 56.91 across the 7 datasets. We also find that Qwen-2.5VL experiences the greatest gain, having an average accuracy difference of +17.46 from the base constrained decoding method to nlg2choiceopen. Comparatively, we see that Intern3-VL performs the worst across architectures, having the overall smallest increase when using nlg2choiceopen (+6.87) and the lowest performance (24.00). However, we do still see improvements over all benchmarks.
@@
A natural question is to ask how the classifications qualitatively change under nlg2choice. For an example, take Fig. 1 where the ground truth label is "Ivory Gull": under constrained decoding the model responds with "Crested Auklet" whereas a higher quality misclassificaiton would be "Herring Gull", "California Gull", …, ie. a class within the same genus as Ivory Gull. In Tab. 3 we qualitatively gauge how misclassifications change when using nlg2choice by measuring its genus-level performance. We construct a species to genus mapping on the CUB200 dataset manually, then map the species prediction in previous results to genus. Simply put, we find that nlg2choice makes more reasonable errors, switching to nlg2choice incurs a genus-level performance improvement of +16.75, +23.85, +23.59 for Qwen-2.5VL, Llama-3.2V-11B, and Intern3VL-8B.
Figure 3: Difference in accuracies on the question-level between nlg2choice and constrained decoding. Each distribution is made up of the performance difference when switching to nlg2choice over 15 questions. nlg2choice uses an open-ended prompt whereas constrained decoding uses an instruction to increase brevity ("... Answer with species only.").
@@
We find that for the proposed settings, switching from constrained decoding with an instruction to increase brevity ("... Answer with species only.") to nlg2choiceopen comes with an average $\mu=\mathbf{22.06},\mathbf{7.38},\mathbf{7.72}$ percentage increase in accuracy for Qwen-2.5VL, Llama-3.2V-11B, and Intern3VL-8B, respectively. Additionally, we find that these increases are within the bounds of statistical significance ($\sigma=\mathbf{1.92},\mathbf{1.69},\mathbf{2.95}$).
@@
We see that for every template, switching from constrained decoding to nlg2choice improves performance. The most stark setting is the most open-ended prompt, where we see an average increase of +68.97, +38.33, and +24.34 for Qwen-2.5VL, Llama-3.2V, and Intern3-VL, respectively. Surprisingly, explicit choice instruction shows the weakest improvement for Qwen-2.5VL and Intern3-VL (+2.82 and +0.81) but strong improvement for Llama-3.2V (+27.19). Otherwise, keeping in line with previous results, we find that Qwen-2.5VL is the greatest benefactor of performance.
@@
... "Let’s think step by step." [16]
Qwen-2.5VL 54.93 (-8.43) 70.16 (-7.87)
@@
it into steps." [30]
Qwen-2.5VL 54.69 (-8.67) 59.56 (-18.47)
@@
In general, we see that these methods do not consistently improve performance, resulting in an average decrease of -9.75, -15.18, and -2.36 across Qwen-2.5VL, Llama-3.2V, and Intern3-VL, respectively. The one exception to this rule is the CUB200 dataset with Intern3-VL, where we see an average increase of +1.01.
@@ 5.2 Zero-Shot Retrieval Performance
In Tab. 6, we report the performance of various models when viewing FGVC datasets as a one-vs-rest retrieval task. Like the classification setting, we find that when compared to exhaustive “Yes/No” questioning, nlg2choice generally improves with the exception of the Stanford Cars [17], in which performance across all models decreases. We again find Qwen-2.5VL to receive the greatest improvement, with +8.16 average mAP increase across the 5 datasets. Likewise, we find Intern3-VL receives the least improvement, with an average increase of +4.43.
@@ 5.3 Answer Extraction Performance
xFinder [47]
Qwen-1.5-0.5B [37] 92.05 88.67 71.76 86.27
Llama-3-8B-Instruct [12] 80.13 80.67 75.29 79.27
@@ Appendix B Zero-Shot 4-Way VQA Performance
Following recent work [36], we evalute nlg2choice on popular fine-graiend classification datasets as a 4-way VQA task. iNat-Animal and iNat-Plant are the rows of iNat21 [39] when subset to rows which whose kingdom matches ”Animalia” and ”Plantae.” ImgNet-Animal and ImgNet-Artifact are the rows of ImageNet-1K [32] subset to rows which have an anceter of ”Animal” and ”Artifact” in the WordNet hierarchy [26]. The choices for each example are from the top 3 SigLIP [48] predictions and the correct class. For the nlg2choice prompt we use the same prompt as the previous work. We report the open-ended question performance in Tab. 8 where we find that it slightly underperforms or has not effect on the lettering approach. Specifically, we find that accuracy changes by -0.03, +0.58, and +0.30 on average for Qwen-2.5VL, Llama-3.2V, and Intern3VL, respectively.
Decreasing the choice set improves performance. We are also able to compare the 4-way VQA setting of CUB200 directly to the many-way prediction displayed in Tab. 2. We see that subsetting the full choice set to four choices incurs a performance improvement of +4.07, +19.20, and +28.83 for Qwen-2.5VL, Llama-3.2V, and Intern3VL, respectively.

### v2

@@ 1 Introduction
Zero-shot visual classification has been revitalized by the introduction of Multimodal Large Language Models (MLLMs) [llava, instructblip, gpt4v, llava_next], where one can formulate classification datasets as Visual Question Answering (VQA) tasks with Multiple Choice Questions (MCQs) and a common prompt. However, fine-grained visual classification (FGVC) poses a barrier to traditional MCQ problems due to two issues: (1) a vastly higher number of choices (often hundreds to thousands) and (2) the niche concepts that each choice represents. Recent evaluations have found that state-of-the-art instruction-tuned MLLMs (e.g. LLaVA 1.5 [llava, llava_next], InstructBLIP [instructblip], GPT-4V [gpt4v]) suffer dramatic drops in accuracy when asked to name fine-grained categories. For instance, LLaVA-1.5 (13B) achieves near-perfect accuracy on coarse categories (like “bird” vs “not bird”), but only  1–2% accuracy on fine-grained species labels in iNaturalist [finer].
Figure 2: Seemingly innocuous instruction changes greatly affect MLLM performance. When testing on the gull subset of CUB200[cub_dataset], replacing one word can change the predictions of a model (left). When rewriting an instruction in a semantically equivalent way, performance varies substantially for Qwen-2.5VL (right). Each distribution comprises 15 semantically equivalent ways of writing the same instruction. These cases highlight the need to test LLMs over various equivalent inputs to ensure robustness.
Using LLMs to extract choices from pre-existing text appears to be a potential solution to this problem. Constraining LLM’s output format to ease extraction can simplify scoring but might hinder the model’s reasoning [wang2025slot, tam2024let, wang2024look], whereas allowing free-form explanations improves reasoning at the cost of harder extraction [molfese2025right, tam2024let]. Many approaches prompt a strong LLM [chiang2023vicuna, li2023alpacaeval] such as GPT4 [achiam2023gpt] to perform second-stage answer extraction. However, the works that propose LLMs as choice extractor are either focused only on language-only tasks [lyu2024beyond, wang2024my] or do not extend beyond simple lettering or numbering schemes [mmbench]. Furthermore, no significant treatment of retrieval nor the robustness of these results under input prompt variation has been considered.
@@ 2 Related Work
The earliest works first prompted an LLM directly for a response then did regex parsing to extract choices [wang2022self], but it was shown that direct open-ended querying often underperforms on fine-grained tasks [finer]. While generating the full probability the text of each choice sequence is as an option, it remains difficult to implement in practice due to its computational cost [robinson2023leveraginglargelanguagemodels]. The most dominant choice probability generation rank the log probabilities assigned to the option IDs (e.g. ”A/B/C/D”) from the first token prediction of the model. This approach is widely adopted in many different benchmarks and LLM evaluation studies [hendrycks2020measuring, wang2024mmlu, srivastava2022beyond, liang2022holistic, santurkar2023whose].
Parallel to this, constrained decoding techniques seek to guarantee consistent outputs by guiding the generation process during sampling by constricting the output vocabulary [beurer2024guiding, liu2024we]. There are many open-source tools that implement this approach [dong2024xgrammar, GuidanceAI2023, Gerganov2023, outlines] and while these methods ensure model responses lie within the choice set, recent evaluations have shown that they decrease performance [wang2025slot, tam2024let]. Our work sits in opposition to these recent results, where we posit that constrained decoding is advantageous in the fine-grained setting where previous approaches like first token prediction fail.
@@
Recent works have shown a mismatch difference between text-based outputs and token-probability metrics of LLMs on language-only tasks, revealing that a model’s first-token likelihood often fails to correspond with its generated answer [lyu2024beyond, wang2024my, wang2024look]. Evaluating a model by reading its full text answer instead of only the first-token or forced letter has been shown to improve robustness, reinforcing the value of answer extraction methods [wang2024look]. Many works use large proprietary models like GPT-4 [achiam2023gpt] to extract answers from free-form text [mmbench, chiang2023vicuna, li2023alpacaeval].
Some recent methods have addressed the problem of answer extraction directly. xFinder [yu2024xfinder] propose a lightweight model fine-tuned specifically to map an LLM’s free-form answer to the correct choice, reporting that their method outperforms a GPT-4 based extractor struggled in terms of extraction accuracy. Similarly, SLOT [wang2025slot] proposes a model-agnostic, lightweight “Structured LLM Output Transformer” that is fine-tuned to post-process any LLM’s free-form text into JSON that exactly matches an arbitrary schema, eliminating the need to modify or constrain the generating model itself. xVerify [xVerify] is an efficient answer verifier for reasoning model evaluations that extracts final answers from complex, multi-step LLM outputs and performs robust equivalence checking—trained on the VAR dataset of multi-model, multi-round–annotated question–answer pairs and achieving over 95% accuracy and F1, demonstrating strong generalizability. Contrary to these approaches the answer extraction method proposed in this work requires no additional training. We argue that answer extraction is not a critical bottleneck when using constrained decoding for fine-grained visual classification and show that LLMs perform decent answer extraction out-of-the-box on popular benchmarks.
@@ 3.1 Simulating User Variation
Where the words ”species” and ”bird” are updated to suite the benchmark dataset. Given [templatei], we use o3-high [o3] to generate 14 additional prompt variations:
@@
Table 1: Number of forward passes needed for different methods in the second-stage answer extraction of nlg2choice. ”Full Prob” refers to the probability of a full class name $p(\texttt{[cname]})$ under the LLM, where [cname] is a common name of a class. “Yes/No” refers to the common $|\mathcal{X}|\cdot|\mathcal{Y}|$ method of calculating probability over all the classes $y_{i}\in\mathcal{Y}$. “Truncated Prob” refers to the process outlined in Sec. $3.$ The percentages in the “Yes/No” and ”Truncated Prob” columns refer to the throughput increase over caculating the full sequence probabilities ”Full Prob.” Every setting for truncation is faster than “Yes/No” except for one benchmark - NABirds [nabirds_dataset]. The fastest method for each benchmark is bolded.
@@ 4 Experiments
We conduct experiments over a benchmark of 7 FGVC datasets: CUB [cub_dataset] (200 classes), Flowers 102 [flowers_dataset] (102 classes), Stanford Cars [cars_dataset] (196 classes), FGVC Aircrafts [aircrafts_dataset] (100 classes) and Food101 [food_dataset] (101 classes), NABirds [nabirds_dataset] (555 classes), and iNaturalist-Birds [inat_dataset] (1486 classes). We choose these benchmarks because they are well-known for fine-grained visual classification and are traditionally difficult for zero-shot visual models to solve. iNaturalist-Birds is the validation split of the iNat2021 dataset where the the ”class” field of each row is equal to ”Aves,” resulting in a test size of $14960$. For each dataset we use the full test set and the training split remains unused.
Evaluation Metrics   For the classification setting we use accuracy averaged over 15 semantically equivalent prompts, whereas for the retrieval setting we use Mean Average Precision (mAP). Choice lists are the names of classes themselves rather than letters or numbers. As compared to previous works [tan2025vision], we perform the full-way class prediction rather than subsetting the choices to a 4-way prediction. However, we provide an experiment evaluating this setting in Appendix B.
@@
Implementation Details   Constrained decoding is implemented using the Outlines [^2] https://github.com/dottxt-ai/outlines library [outlines] and classes for each setting are detailed in Appendix C. We start by generating 15 prompts described in Sec. 3.1, then replacing the key features like domain (”birds,” ”flowers,” etc.) and class types (”species,” ”variant,” etc.) for each dataset. This allows us to compare directly between questions for a given model or dataset, and we also test the performance of appending class lists and common CoT methods. For our experiments, we test only open-source models. Specifically, we test the Qwen-2.5VL [qwen25vl], Llama-3.2 [llama3], and InternVL3 [internvl3] architectures. For hyperparameters, we use the recommended generation parameters as defined by the HuggingFace card and take max new tokens to be 512. Experiments were ran in a heterogeneous environment of single-GPU nodes. All nodes required a GPU with at least 23GB VRAM, 16 CPU cores, and 64GB RAM.
@@ 5.1 Zero-Shot Classification Performance
In Tab. 2 we report the performance of nlg2choice across various architectures and datasets. All architectures enjoyed a bump in performance across individual datasets and on average (+16.97, +8.49, +6.87 across Qwen-2.5VL, Llama-3.2V, and Intern3), but model-specific differences can be clearly seen. Likewise, we find that removing explicit answer formatting from the model (changing from “nlg2choice” and “nlg2choiceopen”). We find the Aves subset of iNaturalist [inat_dataset] to be the hardest among all the benchmarks, resulting in the smallest increase in performance across all models.
For model-specific differences, we find that Qwen-2.5VL performs the best on average using nlg2choice, having an average accuracy of 56.91 across the 7 datasets. We also find that Qwen-2.5VL experiences the greatest gain, having an average accuracy difference of +16.97 from the base constrained decoding method to nlg2choiceopen. Comparatively, we see that Intern3-VL performs the worst across architectures, having the overall smallest increase when using nlg2choiceopen (+6.87) and the lowest performance (24.00). However, we do still see improvements over all benchmarks.
@@
A natural question is to ask how the classifications qualitatively change under nlg2choice. For an example, take Fig. 1 where the ground truth label is "Ivory Gull": under constrained decoding the model responds with "Crested Auklet" whereas a higher quality misclassificaiton would be "Herring Gull", "California Gull", …, ie. a class within the same genus as Ivory Gull. In Tab. 3 we qualitatively gauge how misclassifications change when using nlg2choice by measuring its genus-level performance. We construct a species to genus mapping on the CUB200 dataset manually, then map the species prediction in previous results to genus. Simply put, we find that nlg2choice makes more reasonable errors, switching to nlg2choice incurs a genus-level performance improvement of +17.91, +23.85, +23.59 for Qwen-2.5VL, Llama-3.2V-11B, and Intern3VL-8B.
Figure 3: Difference in accuracies on the question-level between nlg2choice and constrained decoding on CUB200. Each distribution is made up of the performance difference when switching to nlg2choice over 15 questions. nlg2choice uses an open-ended prompt whereas constrained decoding uses an instruction to increase brevity ("... Answer with species only.").
@@
We find that for the proposed settings, switching from constrained decoding with an instruction to increase brevity ("... Answer with species only.") to nlg2choiceopen comes with an average $\mu=\mathbf{18.96},\mathbf{7.38},\mathbf{7.72}$ percentage increase in accuracy for Qwen-2.5VL, Llama-3.2V-11B, and Intern3VL-8B, respectively. Additionally, we find that these increases are within the bounds of statistical significance ($\sigma=\mathbf{2.01},\mathbf{1.69},\mathbf{2.95}$).
@@
We see that for every template, switching from constrained decoding to nlg2choice improves performance. The most stark setting is the most open-ended prompt, where we see an average increase of +67.23, +38.33, and +24.34 for Qwen-2.5VL, Llama-3.2V, and Intern3-VL, respectively. Surprisingly, explicit choice instruction shows the weakest improvement for Qwen-2.5VL and Intern3-VL (+2.82 and +0.81) but strong improvement for Llama-3.2V (+27.19). Otherwise, keeping in line with previous results, we find that Qwen-2.5VL is the greatest benefactor of performance.
@@
... "Let’s think step by step." [kojima2022large]
Qwen-2.5VL 54.93 (-4.95) 70.16 (-7.87)
@@
it into steps." [reynolds2021prompt]
Qwen-2.5VL 54.69 (-5.19) 59.56 (-18.47)
@@
In general, we see that these methods do not consistently improve performance, resulting in an average decrease of -8.01, -15.18, and -2.36 across Qwen-2.5VL, Llama-3.2V, and Intern3-VL, respectively. The one exception to this rule is the CUB200 dataset with Intern3-VL, where we see an average increase of +1.01.
@@ 5.2 Zero-Shot Retrieval Performance
In Tab. 6, we report the performance of various models when viewing FGVC datasets as a one-vs-rest retrieval task. Like the classification setting, we find that when compared to exhaustive “Yes/No” questioning, nlg2choice generally improves with the exception of the Stanford Cars [cars_dataset], in which performance across all models decreases. We again find Qwen-2.5VL to receive the greatest improvement, with +8.16 average mAP increase across the 5 datasets. Likewise, we find Intern3-VL receives the least improvement, with an average increase of +4.43.
@@ 5.3 Answer Extraction Performance
xFinder [yu2024xfinder]
Qwen-1.5-0.5B [qwen1.5] 92.05 88.67 71.76 86.27
Llama-3-8B-Instruct [llama3] 80.13 80.67 75.29 79.27
@@ Appendix B Zero-Shot 4-Way VQA Performance
Following recent work [tan2025vision], we evalute nlg2choice on popular fine-graiend classification datasets as a 4-way VQA task. iNat-Animal and iNat-Plant are the rows of iNat21 [inat_dataset] when subset to rows which whose kingdom matches ”Animalia” and ”Plantae.” ImgNet-Animal and ImgNet-Artifact are the rows of ImageNet-1K [imagenet_dataset] subset to rows which have an anceter of ”Animal” and ”Artifact” in the WordNet hierarchy [wordnet_dataset]. The choices for each example are from the top 3 SigLIP [zhai2023sigmoid] predictions and the correct class. For the nlg2choice prompt we use the same prompt as the previous work. We report the open-ended question performance in Tab. 8 where we find that it slightly underperforms or has not effect on the lettering approach. Specifically, we find that accuracy changes by -0.03, +0.58, and +0.30 on average for Qwen-2.5VL, Llama-3.2V, and Intern3VL, respectively.
Decreasing the choice set improves performance. We are also able to compare the 4-way VQA setting of CUB200 directly to the many-way prediction displayed in Tab. 2. We see that subsetting the full choice set to four choices incurs a performance improvement of +7.55, +19.20, and +28.83 for Qwen-2.5VL, Llama-3.2V, and Intern3VL, respectively.